Circular No. 5713

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

α ORIONIS

A. P. Hatzes and W. D. Cochran, University of Texas at Austin, report: “Precise relative radial velocities of α Ori obtained at McDonald Observatory, using an iodine absorption cell, indicate that this star has exhibited a slow decline in its radial velocity by ∼ 6 km/s from 1991 April to 1992 November. Between 1992 Nov. 12 and 1993 Feb. 2, the radial velocity underwent a rapid increase of ∼ 3 km/s, coinciding with the decrease in brightness for the star reported by Guinan et al. on IAUC 5708.”

PERIODIC COMET SCHAUMASSE (1992x)

Extension to ephemeris on IAUC 5666:

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 4 5^{h}18.^{m}38 +45^{o}08.^{′}1 0.556 1.202 98.^{o}0 54.^{o}8  8.4
9 5 43.24 +46 21.7 0.561 1.204  97.7 54.8  8.5
14 6 10.44 +47 11.2 0.568 1.209  97.7 54.6  8.5
19 6 39.42 +47 32.3 0.577 1.218  98.0 54.0  8.6
24 7 09.43 +47 21.8 0.589 1.231  98.5 53.2  8.7
29 7 39.56 +46 38.9 0.603 1.246  99.2 52.3  8.8
Apr. 3 8 08.95 +45 24.9 0.621 1.265 100.0 51.1  9.0
8 8 36.90 +43 43.3 0.642 1.286 100.8 49.9  9.2
13 9 02.96 +41 38.7 0.667 1.310 101.5 48.6  9.4
18 9 26.93 +39 16.5 0.696 1.337 102.2 47.2  9.6
23 9 48.82 +36 42.1 0.730 1.366 102.7 45.9  9.8
28 10 08.76 +34 00.4 0.767 1.397 103.0 44.6 10.1
May 3 10 26.93 +31 15.6 0.809 1.429 103.2 43.4 10.4

Total visual magnitude estimates: Feb. 10.15 UT, 8.5 (A. Hale, Las Cruces, NM, 0.41-m reflector); 15.92, 8.4 (J.-C. Merlin, Le Creusot, France, 0.40-m reflector); 19.87, 8.4 (H. Lüthen, Hamburg, Germany, 0.20-m reflector); 24.14, 9.3 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

PERIODIC COMET SWIFT-TUTTLE (1992t) Total visual magnitude estimates: Feb. 17.69 UT, 9.1 (A. Jones, Nel-

son, New Zealand, 0.32-m reflector); 18.69, 9.4 (Jones); 19.81, 9.3 (D. A. J. Seargent, The Entrance, N.S.W., Australia, 25×100 binoculars); 23.75, 9.5 (P. Camilleri, Cobram, Vic., Australia, 20×80 binoculars).

1993 February 24 Daniel W. E. Green